3. When data are discrete, we can construct a frequency distribution in which each possible value of the variable forms a class. The following table and histogram presents the results of a hypothetical survey in which 1000 adult women were asked how many children they had.
Number of children Frequency
0 435
1 175
2 222
3 112
4 38
5 9
6 7
7 0
8 2
1
Expert's answer
2019-09-12T09:07:48-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
\hline
 \begin{matrix}
 Class \\
 Limits
\end{matrix} & \begin{matrix}
 Class \\
 Boundaries
\end{matrix} &\begin{matrix}
 Frequency \\
\end{matrix} &\begin{matrix}
 Relative \\
 Frequency
\end{matrix} \\ \hline
-0.5\div0.5 & 0 & 435 &0.435 \\ \hline
0.5\div1.5 & 1 & 175 &0.175 \\ \hline
1.5\div2.5 & 2 & 222 &0.222 \\ \hline
2.5\div3.5 & 3 & 112 &0.112 \\ \hline
3.5\div4.5 & 4 & 38 &0.038 \\ \hline
4.5\div5.5 & 5 & 9 &0.009 \\ \hline
 5.5\div6.5 & 6 & 7 & 0.007\\
\hline
 6.5\div7.5 & 7 & 0 & 0\\
\hline
 7.5\div8.5 & 8 & 2 & 0.002\\
\hline
\end{array}"